How do you find the slope of a line parralel to the graph of each equation? Question: #1 y=-4. #2 -8x+6y=4 and #3 0.5x-6y=4.. I even go to coach class.. Still no help.

Found 2 solutions by rfer, solver91311:
Answer by rfer(16322)

put in y=mx+b form
2)
-8x+6y=4
6y=8x+4
y=8/6x+4/6
y=4/3x+2/3
Slope m=4/3

Answer by solver91311(24713)





In other words, parallel lines have equal slopes. Just determine the slope of your given line and you have the slope of any parallel.

Hint for #1. This equation describes a set of ordered pairs where the coordinate can be any real number and the coordinate is always -4]


John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism